OpenStudy (anonymous):

This problem refers to a right triangle ABC with C=90∘. If a=10 and b=24, find c. I'm stuck? How do I go about solving this problem.

OpenStudy (noseboy908):

In this situation, we can use the Pythagorean theorem to solve for c. Do you know what that is?

OpenStudy (mathmale):

You are given an extremely important clue here: this triangle is a right triangle. That means that one of the angles is 90 degrees, and the other two angles are between 0 and 90 degrees. Further more, you are told that the side (c) OPPOSITE the right angle (C) is to be found, given that the two shorter sides of the triangle are a = 10 and b = 24. Why not draw this information? Then think of tools you've used before in solving right triangles. Look what noseboy908 is suggesting.

OpenStudy (anonymous):

a^2+b^2=c^2 right?

OpenStudy (noseboy908):

Correct! Now, what two parts are we given?

OpenStudy (anonymous):

ok I squared, 10 and 24 and I got 676=c^2

OpenStudy (noseboy908):

Correct, so what's c?

OpenStudy (anonymous):

I'm assuming I sqrt(676)=c right?

OpenStudy (noseboy908):

Yep

OpenStudy (mathmale):

...with one caveat: we take and use only the positive square root of 676. (Why?) Can Sqrt(676) be reduced? Why or why not?

OpenStudy (anonymous):

c=26?

OpenStudy (noseboy908):

Correct! That's your final answer.

OpenStudy (anonymous):

thank you!

OpenStudy (noseboy908):

*fancily bows* Thou art most welcome.